OpenStudy (anonymous):

To find the quotient , multiply numerator and denominator by 4 + 5i

OpenStudy (anonymous):

if you do that then (4 + 5i) is the complex conjugate of the denominator (4 - 5i) which will provide you a common denominator of \[(4+5i)(4-5i)=4^2-5^2i^2=16-25(-1)=14+25=39\]